OpenStudy (anonymous):

(1-cos²θ)(1+cos²θ)=2sin²θ-sin⁴θ. Need to show this is an identity.

OpenStudy (anonymous):

\[(1-\cos^2\alpha)(1+\cos^2\alpha) = \sin^2\alpha[1+(1-\sin^2\alpha)]\]